OpenStudy (anonymous):

To help his sister, Lydia, pay for college, Jason offered her a loan at a simple annual interest rate of 2%. At the end of 4 years, Lydia paid off the loan, and Jason received $80 in interest. How much money did Jason loan his sister?

jimthompson5910 (jim_thompson5910):

Use the formula I = P*r*t I = interest P = principal (ie amount loaned) r = interest rate t = time in years

OpenStudy (anonymous):

how do u get the answer.

jimthompson5910 (jim_thompson5910):

In your case, I = 80 P = unknown (leave it as P for now) r = 0.02 t = 4

jimthompson5910 (jim_thompson5910):

0.02 = 2% since 2% = 2/100 = 0.02

jimthompson5910 (jim_thompson5910):

plug those values in and solve for P

jimthompson5910 (jim_thompson5910):

I = P*r*t 80 = P*0.02*4 P = ???

jimthompson5910 (jim_thompson5910):

what is 0.02*4 equal to

OpenStudy (anonymous):

0.08

jimthompson5910 (jim_thompson5910):

we now have 80 = P*0.08

jimthompson5910 (jim_thompson5910):

we isolate P by undoing that multiplication of 0.08

jimthompson5910 (jim_thompson5910):

so we divide both sides by 0.08 to isolate P

OpenStudy (anonymous):

how do u do that

jimthompson5910 (jim_thompson5910):

80/0.08 = ???

OpenStudy (anonymous):

omg thanks

OpenStudy (anonymous):

When Janette was 5 years old, her grandmother deposited $2500 into a savings account for her. The account pays a simple annual interest rate of 4%. What will the balance be when Janette turns 15 years old?

jimthompson5910 (jim_thompson5910):

same idea, just slightly different

jimthompson5910 (jim_thompson5910):

you use I = P*r*t in this case, P = 2500, r = 0.04, t = 15 calculate the interest earned and add that to 2500